# Asymptotic Behavior of Rupture Solutions for the Elliptic MEMS Equation with Hénon-Type Term

### v1

@@ 1 Introduction
In this sction, we give the arbitrary order asymptotic expansion of the case
@@
Assume $u\in C^{2}(B\backslash\{0\})$ is a positive rupture solution and exists $\epsilon>0$ satisfying $u(x)=u_{s}(|x|)\left(1+O\left(|x|^{\epsilon}\right)\right)$. Define $z(t,\theta)=r^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$,( $\Lambda$ is same as above), $\mathrm{t}=\ln|x|$, then it exists a positive sequence $\left\{\mu_{j}\right\}_{j\geq 1}$ strictly increasing to $\infty$.
When $F=0$, we have
$\mu_{1}=\begin{cases}\sigma_{1}^{(1)},&-2<\alpha<\delta^{(1)},\\[2.0pt] \sigma_{1}^{(k)},&\delta^{(k-1)}<\alpha<\delta^{(k)},\quad k=2,3,\cdots.\end{cases}$
When $F\neq 0$.
${u_{1}}=\left\{\begin{array}[]{lll}{\sigma_{1}^{(1)}}&-2<\alpha<\delta^{(1)}~and~\alpha<2p-(p+1)\sigma_{1}^{(1)}\\[2.0pt] {\sigma_{1}^{(k)}}&\delta^{(k-1)}<\alpha<\delta^{(k)}~and~\alpha<2p-(p+1)\sigma_{1}^{(k)}\\ \displaystyle\frac{2p-\alpha}{p+1}&\text{if $\delta^{(k-1)}<\alpha<\delta^{(k)}$ and $\alpha\geq 2p-(p+1)\sigma_{1}^{(k)}$,}\end{array}\right$.
For any positive interger $n\gg 1$ and $(t,\theta)\in(-\infty,-1)\times S^{N-1}$
$z(t,\theta)=\sum_{j=1}^{n}\sum_{l=0}^{j-1}c_{jl}(\theta)t^{l}e^{\mu_{j}t}+O\left(t^{n}e^{\mu_{n+1}t}\right)$
[…]
$\Delta_{S^{N-1}}Q(\theta)=\lambda_{i}Q(\theta)$
proof. It’s easy to see $z(t,\theta)=O(e^{\epsilon t})$ as $t\to$-$\infty$, and it satisfies the equation
@@
$\Delta_{S^{N-1}}Q=\lambda Q$
$\lambda_{k}=k(N-2+k)$ and
$m_{k}=\frac{(N-2+2k)(N-3+k)!}{k!(N-2)!}.$
We define $\{Q_{1}^{k}(\theta),\ldots,Q_{m_{k}}^{k}(\theta)\}$ is a set of orthonormal bases corresponding to the feature space of $\lambda_{k}$
@@
$\sigma_{1}^{(k_{0}+1)}>\sigma_{1}^{(k_{0})}>\sigma_{1}^{(1)}>0,\quad\sigma_{2}^{(k_{0}+1)}<\sigma_{2}^{(k_{0})}<\sigma_{2}^{(1)}<0.$
where $2<\alpha<\delta^{(1)}$ ($k=k_{0}\geq 2$).
$\sigma_{1}^{(k_{0}+1)}>\sigma_{1}^{(k_{0})}>0>\sigma_{1}^{(k_{0}-1)},\quad\sigma_{2}^{(k_{0}+1)}<\sigma_{2}^{(k_{0})}<\sigma_{2}^{(1)}<0.$
where $\delta^{(k_{0}-1)}<\alpha<\delta^{(k_{0})}$ ($k=k_{0}\geq 2$)
Remark: In fact,It will be possibile that $\sigma_{1}^{(k)}$ and $\sigma_{2}^{(k)}$ are imaginary number when k is samll enough, but it will be similar with the proof below because such $\sigma^{(k)}$ is finite and $\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)<0$.
@@
For $N\geq 2$, $p>0$. Assume $u$ is a positive rupture solution of (1) satisfying $u(x)=u_{s}(|x|)\left(1+O\left(|x|^{\epsilon}\right)\right)$ with
$\left\{\begin{array}[]{lll}{0<\epsilon\leq\sigma_{1}^{(1)},\qquad\text{as }-2<\alpha<\delta^{(1)}~~and~~\alpha<2p-(p+1)\sigma_{1}^{(1)}}\\ {0<\epsilon\leq\sigma_{1}^{(k)},\qquad\text{as }\delta^{(k-1)}<\alpha<\delta^{(k)}~~and~~\alpha<2p-(p+1)\sigma_{1}^{(k)}}\\ {0<\epsilon\leq\frac{2p-\alpha}{p+1},\qquad\text{as }\delta^{(k-1)}<\alpha<\delta^{(k)}~~and~~\alpha\geq 2p-(p+1)\sigma_{1}^{(k)}}\end{array}\right$.
If we define $z(t,\theta)=|x|^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$ ($t=\ln r$), then $z(t,\theta)=\mathcal{O}(e^{\varepsilon t})$ for t $\in(-\infty,1)$ , and for the case $F\neq 0$,
[…]
For the case $F=0$,
$\max_{S^{N-1}}|z(t,\theta)|\leq\left\{\begin{array}[]{ll}{\sigma_{1}^{(1)}}&\text{if }-2<\alpha<\delta^{(1)},\\ {\sigma_{1}^{(k)}}&\text{if }\delta^{(k-1)}<\alpha<\delta^{(k)}.\end{array}\right$.
@@
$\max_{S^{N-1}}|w(t,\theta)|\leq\left\{\begin{array}[]{ll}\sigma_{1}^{(1)},&\text{if }-2<\alpha<\delta^{(1)},\\[2.84526pt] \sigma_{1}^{(k)},&\text{if }\delta^{(k-1)}<\alpha<\delta^{(k)}.\end{array}\right$.
we focus on the case $\delta^{(k-1)}<\alpha<\delta^{(k)}$, the remaining case is similiar and eaiser.
@@
$\|w\|_{L^{2}(\mathbb{S}^{N-1})}^{2}=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{(k)}(t)\right)^{2},\quad|f(z)-f(\overline{z})|=\sum_{k=1}^{\infty}\sum_{j=0}^{m_{k}}\left(g_{j}^{(k)}(t)\right)^{2}$
@@
$w_{j}^{k}(t)=O\left(e^{\sigma_{1}^{(k)}(t-T)}\right)+B_{j,1}^{k}\int_{t}^{T}e^{\sigma_{1}^{(k)}(t-s)}g_{j}^{k}(s)\,ds-B_{j,2}^{k}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{j}^{k}(s)\,ds.$ (12)
@@
$w_{j}^{k}(t)=$ $A_{j,1}^{k}e^{\sigma_{1}^{(k)}t}+A_{j,2}^{k}e^{\sigma_{2}^{(k)}t}-B_{j,1}^{k}\int_{-\infty}^{t}e^{\sigma_{1}^{(k)}(t-s)}g_{j}^{k}(s)ds-B_{j,2}^{k}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{j}^{k}(s)ds$
@@
$\|f^{\prime}(\xi)w\|_{L^{2}(S^{N-1})}^{2}=O\left(e^{2\varepsilon t}\right)\|w\|_{L^{2}(S^{N-1})}^{2}=O\left(e^{2\varepsilon t}\right)\left[\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}+\sum_{k=1}^{k_{0}-1}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}\right]$ (3.1)
@@
${}^{2}\leq$ $\;Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+Ce^{4\varepsilon t}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{4\varepsilon s}\,ds$
@@
${}^{2}\leq$ $Ce^{(2\sigma_{1}^{(2)}-\delta)(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(2)}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds$ (16)
@@
${}^{2}\leq$ $Ce^{4\varepsilon t}+C\int_{t}^{T}e^{(2\sigma_{1}^{(2)}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds$
@@
$\begin{gathered}{}^{2}\leq Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{6\varepsilon s}\,ds+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds\\ \quad+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{6\varepsilon s}\,ds+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.\end{gathered}$ (28)
@@
$\begin{gathered}{}^{2}\leq Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{8\varepsilon s}\,ds+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds\\ \quad+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{8\varepsilon s}\,ds+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.\end{gathered}$
@@
This completes the proof of the lemma.
@@
$|\bar{z}(t)|\leq\begin{cases}Ce^{\min\left\{2\sigma_{1}^{(1)},\,\frac{2p-\alpha}{p+1}\right\}t},&\text{if }-2<\alpha<\delta^{(1)},\\[2.84526pt] Ce^{\min\left\{2\sigma_{1}^{(k)},\,\frac{2p-\alpha}{p+1}\right\}t},&\text{if }\delta^{(k-1)}<\alpha<\delta^{(k)}.\end{cases}$
For $F\neq 0$,
$|\bar{z}(t)|\leq\begin{cases}Ce^{2\sigma_{1}^{(1)}t},&\text{if }-2<\alpha<\delta^{(1)},\\[2.84526pt] Ce^{2\sigma_{1}^{(k)}t},&\text{if }\delta^{(k-1)}<\alpha<\delta^{(k)}.\end{cases}$
For $F=0$. $C>0$ is a constant independent of $t$.
Proof. We only prove the case $F\neq 0$ with $\delta^{(k-1)}<\alpha<\delta^{(k)}$, the other cases are similar and eaiser. we only need to prove the estimate for $t$ sufficiently close to -$\infty$ and the other part can be gotten easily by the continuity of $\bar{z}(t)$,
@@
where $\varepsilon>0$ is as in Lemma 2.3.
@@
We still consider two cases: (a) $4\varepsilon\geq\frac{2p-\alpha}{p+1}$, (b) $4\varepsilon<\frac{2p-\alpha}{p+1}$. Repeating the similar steps as above then we yield the desired estimate. We now prove Lemma 2.3. From the arguments given above, observe that when $\alpha<(\geq)\,2p-(p+1)\sigma_{1}^{(k)}$, we have $\sigma_{1}^{(k)}<(\geq)\,\frac{2p-\alpha}{p+1}$. Since $z(t,\theta)=w(t,\theta)+\bar{z}(t)$, the conclusion follows immediately.
@@
(i) For the case $F=0$ and $2<\alpha<\delta^{(1)}$, we set
$\rho_{1}=\sigma_{1}^{(1)},\;\rho_{2}=\sigma_{1}^{(2)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m)},\;\ldots$ (29)
For the case $F=0$ and $\delta^{(k-1)}<\alpha<\delta^{(k)}$, we set
$\rho_{1}=\sigma_{1}^{(k)},\;\rho_{2}=\sigma_{1}^{(k+1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(k+m)},\;\ldots$ (30)
(ii) For the case $F\neq 0$ and $2<\alpha<\delta^{(1)}$, if there exists some $j$ such that
$\rho_{1}^{(j)}<\frac{2p-\alpha}{p+1}<\rho_{1}^{(j+1)},$
we set
$\rho_{1}=\sigma_{1}^{(1)},\;\ldots,\;\rho_{j+1}=\frac{2p-\alpha}{p+1},\;\rho_{j+2}=\sigma_{1}^{(j+1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m)},\;\ldots$ (31)
For the case $F\neq 0$ and $2<\alpha<\delta^{(1)}$, if there exists some $j$ such that
we set
$\rho_{1}=\sigma_{1}^{(1)},\;\ldots,\;\rho_{j}=\sigma_{1}^{(j)},\;\rho_{j+1}=\sigma_{1}^{(j+1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m)},\;\ldots$ (32)
For the case $F\neq 0$ and $2<\alpha<\delta^{(1)}$, if
$\frac{2p-\alpha}{p+1}<\sigma_{1}^{(1)},$
we set
$\rho_{1}=\frac{2p-\alpha}{p+1},\;\rho_{2}=\sigma_{1}^{(1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(m-1)},\;\ldots$ (33)
@@
We first prove case (29), and then indicate the proof of the remaining cases.
@@
We now introduce the notion of an index set. Let $\{\rho_{i}\}_{i\geq 1}$ denote the sequence appearing in Lemma 3.5 (i.e., the sequence $\{\rho_{i}\}$ in (29) with multiplicities considered), which is strictly increasing and diverges to $\infty$.
@@
By the definition of $\mathcal{I}_{\widetilde{\rho}}$, it exists that $\widetilde{\rho}_{i}=n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}}$. Thus, by Lemma 3.6,
@@
All remaining steps follow the same pattern as in the proof of (29). Thus Theorem 3.1 is now proved in full.

### v3

@@ 3 Asymptotic behavior for nonradial with asymptotic radial solution
In this section, we give the arbitrary order asymptotic expansion for the asymptotic radial case.
@@
Assume $u\in C^{2}(B\backslash\{0\})$ is a positive rupture solution of (1) and there exists $\epsilon>0$ satisfying $u(x)=\Lambda r^{\frac{\alpha+2}{p+1}}\left(1+O\left(|x|^{\epsilon}\right)\right)$. Let $z(t,\theta)=r^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$, $\Lambda$ is the same as in Theorem 1.2, $t=\ln|x|$, $\theta=\frac{x}{|x|}$, then there exists a positive sequence $\left\{\mu_{j}\right\}_{j\geq 1}$ strictly increasing to $\infty$, for any positive integer $n\gg 1$ and $(t,\theta)\in(-\infty,-1)\times S^{N-1},$
$z(t,\theta)=\sum_{j=1}^{n}\sum_{l=0}^{j-1}c_{jl}(\theta)t^{l}e^{\mu_{j}t}+O\left(t^{n}e^{\mu_{n+1}t}\right),$
where $\sigma_{1}^{(k)}$ and $\delta^{(k)}$ are the same as in Theorem 1.2, $c_{jl}(\theta)=\sum_{i=0}^{m_{jl}}a_{jli}Q_{i}(\theta)$, $a_{jli}$ is a constant, $m_{jl}$ is a integer depending on $N,j,l,p,a$. $Q_{i}(\theta)$ is a linear combination of characteristic functions of
$\Delta_{S^{N-1}}Q(\theta)=\lambda_{i}Q(\theta).$
It’s easy to see that $z(t,\theta)=O(e^{\epsilon t})$ as $t\to$-$\infty$ and
@@
$\Delta_{S^{N-1}}Q=\lambda Q,$
$\lambda_{k}=k(N-2+k)$ and $m_{k}=\frac{(N-2+2k)(N-3+k)!}{k!(N-2)!}$ is the corresponding multiplicity.
We define $\{Q_{1}^{k}(\theta),\ldots,Q_{m_{k}}^{k}(\theta)\}$ to be an orthonormal basis corresponding to the eigenspace of $\lambda_{k}$ .
@@
$\sigma_{1}^{(k)}=-\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)+\frac{1}{2}\sqrt{\left(N-2+2\frac{\alpha+2}{p+1}\right)^{2}+4k(N-2+k)-4(\alpha+2)\left(N-2+\frac{\alpha+2}{p+1}\right)},$
@@
$\sigma_{1}^{(k_{0}+1)}>\sigma_{1}^{(k_{0})}>\sigma_{1}^{(1)}>0,\quad\sigma_{2}^{(k_{0}+1)}<\sigma_{2}^{(k_{0})}<\sigma_{2}^{(1)}<0$
where $2<\alpha<\delta^{(1)}$ ($k_{0}\geq 2$).
$\sigma_{1}^{(k_{0}+1)}>\sigma_{1}^{(k_{0})}>0>\sigma_{1}^{(k_{0}-1)},\quad\sigma_{2}^{(k_{0}+1)}<\sigma_{2}^{(k_{0})}<\sigma_{2}^{(1)}<0$
where $\delta^{(k_{0}-1)}<\alpha<\delta^{(k_{0})}$ ($k_{0}\geq 2$)
Remark. In fact, it will be possible that $\sigma_{1}^{(k)}$ and $\sigma_{2}^{(k)}$ are imaginary numbers when k is sufficiently small. Thew arguments is similar to the proof below because such $\sigma^{(k)}$ is finite and $\frac{1}{2}\left(N-2+2\frac{\alpha+2}{p+1}\right)<0$.
@@
For $N\geq 2$, $p>0$, $k\geq 1$, assume $u$ is a positive rupture solution of (1) satisfying $u(x)=u_{s}(|x|)\left(1+O\left(|x|^{\epsilon}\right)\right)$ with
$\begin{cases}0<\epsilon\leq\sigma_{1}^{(k)},&\text{as }\delta^{(k-1)}<\alpha<\delta^{(k)}~~and~~\alpha<2p-(p+1)\sigma_{1}^{(k)},\\ 0<\epsilon\leq\frac{2p-\alpha}{p+1},&\text{as }\delta^{(k-1)}<\alpha<\delta^{(k)}~~and~~\alpha\geq 2p-(p+1)\sigma_{1}^{(k)}.\end{cases}$
If we define $z(t,\theta)=|x|^{-\frac{\alpha+2}{p+1}}u(x)-\Lambda$, then $z(t,\theta)=\mathcal{O}(e^{\varepsilon t})$ for t $\in(-\infty,1),$ and
$\max_{S^{N-1}}|z(t,\theta)|\leq\begin{cases}\sigma_{1}^{(k)}&\text{as }\delta^{(k-1)}<\alpha<\delta^{(k)}\text{ and }\alpha<2p-(p+1)\sigma_{1}^{(k)},\\ \frac{2p-\alpha}{p+1}&\text{as }\delta^{(k-1)}<\alpha<\delta^{(k)}\text{ and }\alpha\geq 2p-(p+1)\sigma_{1}^{(k)}\end{cases}$
for the case $F\neq 0$.
$\max_{S^{N-1}}|z(t,\theta)|\leq{\sigma_{1}^{(k)}}$
for the case $F=0$.
@@
$\max_{S^{N-1}}|w(t,\theta)|\leq\sigma_{1}^{(k)},$
Proof.
@@
$\|w\|_{L^{2}(\mathbb{S}^{N-1})}^{2}=\sum_{k=1}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{(k)}(t)\right)^{2},\quad|f(z)-f(\overline{z})|=\sum_{k=1}^{\infty}\sum_{j=0}^{m_{k}}\left(g_{j}^{(k)}(t)\right)^{2}.$
@@
$w_{j}^{k}(t)=O\left(e^{\sigma_{1}^{(k)}(t-T)}\right)+B_{j,1}^{k}\int_{t}^{T}e^{\sigma_{1}^{(k)}(t-s)}g_{j}^{k}(s)\,ds-B_{j,2}^{k}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{j}^{k}(s)\,ds.$
@@
$w_{j}^{k}(t)=$ $A_{j,1}^{k}e^{\sigma_{1}^{(k)}t}+A_{j,2}^{k}e^{\sigma_{2}^{(k)}t}-B_{j,1}^{k}\int_{-\infty}^{t}e^{\sigma_{1}^{(k)}(t-s)}g_{j}^{k}(s)ds-B_{j,2}^{k}\int_{-\infty}^{t}e^{\sigma_{2}^{(k)}(t-s)}g_{j}^{k}(s)ds.$
@@
$\|f^{\prime}(\xi)w\|_{L^{2}(S^{N-1})}^{2}=O\left(e^{2\varepsilon t}\right)\|w\|_{L^{2}(S^{N-1})}^{2}=O\left(e^{2\varepsilon t}\right)\left[\sum_{k=k_{0}}^{\infty}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}+\sum_{k=1}^{k_{0}-1}\sum_{j=1}^{m_{k}}\left(w_{j}^{k}(t)\right)^{2}\right].$ (3.1)
@@
${[W(t)]}^{2}\leq$ $\;Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+Ce^{4\varepsilon t}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{4\varepsilon s}\,ds$
@@
${[W(t)]}^{2}\leq$ $Ce^{(2\sigma_{1}^{(2)}-\delta)(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(2)}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds$ (17)
@@
${[W(t)]}^{2}\leq$ $Ce^{4\varepsilon t}+C\int_{t}^{T}e^{(2\sigma_{1}^{(2)}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds$
@@
$\begin{gathered}{[W(t)]}^{2}\leq Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{6\varepsilon s}\,ds+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds\\ \quad+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{6\varepsilon s}\,ds+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.\end{gathered}$ (29)
@@
$\begin{gathered}{[W(t)]}^{2}\leq Ce^{2\sigma_{1}^{(k_{0})}(t-T)}+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{8\varepsilon s}\,ds+C\int_{t}^{T}e^{(2\sigma_{1}^{(k_{0})}-\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds\\ \quad+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{8\varepsilon s}\,ds+C\int_{-\infty}^{t}e^{(2\sigma_{2}^{(k_{0})}+\delta)(t-s)}e^{2\varepsilon s}[W(s)]^{2}\,ds.\end{gathered}$
@@
This completes the proof of the lemma. ∎
@@
$|\bar{z}(t)|\leq Ce^{\min\left\{2\sigma_{1}^{(k)},\,\frac{2p-\alpha}{p+1}\right\}t}$
for $F\neq 0$.
$|\bar{z}(t)|\leq Ce^{2\sigma_{1}^{(k)}t}$
for $F=0$, $C>0$ is a constant independent of $t$.
Proof.
We only prove the case $F\neq 0$, the other cases are similar and easier. we only need to prove the estimate for $t$ sufficiently close to -$\infty$ and the other part can be obtained easily by the continuity of $\bar{z}(t)$.
@@
where $\varepsilon>0$ as in Lemma 3.2.
@@
We still consider two cases: (a) $4\varepsilon\geq\frac{2p-\alpha}{p+1}$, (b) $4\varepsilon<\frac{2p-\alpha}{p+1}$. Repeating the similar steps as above then we yield the desired estimate. ∎
We now prove Lemma 3.2. From the arguments given above, observe that when $\alpha<(\geq)\,2p-(p+1)\sigma_{1}^{(k)}$, we have $\sigma_{1}^{(k)}<(\geq)\,\frac{2p-\alpha}{p+1}$. Since $z(t,\theta)=w(t,\theta)+\bar{z}(t)$, the conclusion follows immediately.
@@
Proof.
(i) For the case $F=0$ and $\delta^{(k-1)}<\alpha<\delta^{(k)}$, we set
$\rho_{1}=\sigma_{1}^{(k)},\;\rho_{2}=\sigma_{1}^{(k+1)},\;\ldots,\;\rho_{m}=\sigma_{1}^{(k+m-1)},\;\ldots$ (30)
(ii) For the case $F\neq 0$ and $\delta^{(k-1)}<\alpha<\delta^{(k)}$, if there exists some $j\geq k$ such that
@@
We prove the subcase that $\delta^{(0)}<\alpha<\delta^{(1)}$ of case (30) at first, and then indicate the proof of the remaining cases.
@@
We now introduce the notion of an index set. Let $\{\rho_{i}\}_{i\geq 1}$ denote the sequence appearing in Lemma 3.5 (i.e. the sequence $\{\rho_{i}\}$ in (30) for $k=1$ with multiplicities considered), which is strictly increasing and diverges to $\infty$.
@@
By the definition of $\mathcal{I}_{\widetilde{\rho}}$, there exists that $\widetilde{\rho}_{i}=n_{1}\rho_{1}+\cdots+n_{r_{1}}\rho_{r_{1}}$. Thus, by Lemma 3.6,
@@
All remaining steps follow the same pattern as in the proof of (30). Thus Theorem 3.1 is now proved in full. ∎